OpenStudy (cometailcane):

The combustion of hydrocarbons produces carbon dioxide and water. The combustion of propane, C3H8, produced 56.2 liters of carbon dioxide at 34.3 degrees C and 103.3 kPa. a. write balanced equation b.determine the moles of carbon dioxide produced, using ideal gas law c.using balanced equation, determine moles of propane used d. determine the mass of propane used

mathslover (mathslover):

C3H8 + 5O2 --> 3CO2 + 4H2O (Balanced Equation) PV = nRT (Ideal Gas Equation) From the above equation, n = (PV)/(RT) You've been provided with : V = 56.2 L ; T = 34.3 Degrees C = 34.3 + 273 K = 307.3 K P = 103.3 * 10^3 Pa Now, use appropriate value of Universal Gas Constant (R) based on the units of the above quantities. ( Reference Link : http://www.csupomona.edu/~lllee/gasconstant.pdf ) You may use the formula then and manipulate the values of n for CO2. Further use The balanced equation to find the moles or propane (Limiting Reagent might come into use).